OpenStudy (anonymous):

Find the Solution of the System of equations: 4x-5y=-52 -4x+3y=44

OpenStudy (thehollywoodstud):

k so lets solve the first equation

OpenStudy (anonymous):

I am horrible with probability. Because it's 'fake' math. But any help would be fantastic.

OpenStudy (thehollywoodstud):

no probability

OpenStudy (solomonzelman):

add the equations

OpenStudy (solomonzelman):

add the equations (do you know what I mean by that?)

OpenStudy (thehollywoodstud):

illl walk her through it

OpenStudy (thehollywoodstud):

combine like terms

OpenStudy (solomonzelman):

there are no like terms here.

OpenStudy (thehollywoodstud):

the x's and the y's

OpenStudy (anonymous):

uhh.. It's getting slightly confusing with you two lol..

OpenStudy (anonymous):

Do you know the elimination method?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 4x~~~~~~~~~~~~~~~~~~-5y~~~~~~~~=~~~~-52 }\) \(\large\color{black}{ \displaystyle -4x~~~~~~~~~~~~~~~~+3y~~~~~~~~=~~~~44 }\) \(\large\color{black}{ \displaystyle \text{______________________________} }\) \(\large\color{black}{ \displaystyle (4x-4x)~~+~~~(-5y+3y)~~~~=~~~~(-52+44) }\)

OpenStudy (solomonzelman):

the x-s cancel out, you see that (right?)

OpenStudy (anonymous):

As you can tell, he added the equations together (using the elimination method)

OpenStudy (anonymous):

Yes, I understand how to eliminate the x's and y's but what I am confused by is what kind of answer Plato is wanting me to give. I've done these problems trillion of different ways and they're always wrong.

OpenStudy (solomonzelman):

it wants an (x,y) coordinate

OpenStudy (anonymous):

Ok, so when you add the equations together, what did you get?

OpenStudy (anonymous):

Yes I see, thank you. I was doing a different method.

OpenStudy (anonymous):

Were you doing the subsitution method?

OpenStudy (anonymous):

Spelled that wrong, whoopsie

OpenStudy (anonymous):

Yeah that's what it told me to do in the tutorial on the site.

OpenStudy (solomonzelman):

I would do a matrix with more variables, but here elimination is the best method.

OpenStudy (anonymous):